OpenStudy (anonymous):

A projectile launched from the ground with an initial velocity 20sqrt5 m/s , just clears two walls each 50m high and seperated by 100m what is the time of passage b/w the walls?

mathslover (mathslover):

first write given thingies : u = 20sqrt5 m/s vertical height = 50 m horizontal height ( between walls) =100m

mathslover (mathslover):

s = ut+1/2 at2 ..

mathslover (mathslover):

hmn wait

mathslover (mathslover):

i got it --> i think

OpenStudy (anonymous):

@mathslover it is the prob frm motion in 2d so resolving vectors is imp

mathslover (mathslover):

like using eqn of motion ..

mathslover (mathslover):

ok first calculate time of flight

mathslover (mathslover):

...... confused........

OpenStudy (dls):

Can you assume the walls to be the highest point of the body and apply the formula of max height?

OpenStudy (anonymous):

it is said it hust cleared so i think we can take it as Hmax

OpenStudy (dls):

Apply the formula and tell me the answer.

OpenStudy (anonymous):

@DLS we need to find what is the time of passage b/w the walls?

OpenStudy (maheshmeghwal9):

 Let the two pillars be covered in t1 & t2 seconds. then we get 3 equations \[50=20\sqrt{5}\sin \theta t_1-\frac{1}{2} g t_1^2.\]............ (1.) eq. \[50 = 20 \sqrt{5} \sin \theta t_2-\frac{1}{2} g t_2^2.\].............(2) eq. \[\color{red}{100=20 \sqrt{5}\cos \theta (t_2-t_1).}\]..............(3.) eq.